OpenStudy (anonymous):

2/(2+n)^1/2 (from zero to infinity) converge or diverge? and how?

OpenStudy (perl):

is this a series

OpenStudy (anonymous):

\[ \sum_1^\infty \frac 1 {n^p} \] diverges if \( p\le1\) and converges if p>1

OpenStudy (anonymous):

Your series is similar to the series I mentioned with p=1/2, so it diverges

OpenStudy (anonymous):

how does it similiar?

OpenStudy (anonymous):

n and 2+n are almost the same near infinity

OpenStudy (anonymous):

Use the limit comparison test

OpenStudy (anonymous):

how

OpenStudy (anonymous):

Do you know the limit comparison test?

OpenStudy (anonymous):

ı guess ı will take limits of both sides right?

OpenStudy (anonymous):

ok thanks a lot

OpenStudy (anonymous):

Read the link above and things will be clearer

OpenStudy (anonymous):

YW

OpenStudy (anonymous):

may ı ask a question? could ı think it like this? : 2. 1/(n+2)^1/2 than say the 1/2 is smaller than 1 so it diverges??

OpenStudy (anonymous):

\[ \lim_{n\to \infty} \frac{2 { (n+2)^{1/2}} } {\frac 1 {n^{1/2}}}=2 >0 \] So your series behaves like \[ \sum_{n=1^{\infty}}\frac 1 {n^{1/2}} \] which is divergent

OpenStudy (anonymous):

Typo \[ \sum_{n=1}^\infty\frac 1 {n^{1/2}} \]